A rectangle is 5cm longer than it is wide. The diagonal of the rectangle is 18cm. What are the
dimensions of the rectangle?

Answer by ikleyn(52810)

.

x^2 + (x+5)^2 = 18^2


x^2 + x^2 + 10x + 25 = 324


2x^2 + 10x + 25 - 324 = 0


2x^2 + 10x - 299 = 0


Solve using the quadratic formula and chose the positive root.